I (30 pts.) A steam power plant with a two-stage turbine and a closed feedwater heater is used to generate power at a local pH1= 3756.1

S1 = 6.9665

H2= 3143.8

H3 = 3690.1

H4 = 2441.8

H5 = 191.85

H6 = 193.8

stuck on H7
Answer #1

from Steam.bababes 9 11859°c) 3 : 120 bar At 120 bar, 659°C 18 20 barn barwy Wy h = 3156 Kolkg $=6.96.65 Kulky- s oolbar Se AAt zobar, 6ooc h = 369011 kJ /kg S2 = 7.702 kJ / kg - K = S4 Atolbar he=191.9 kJ/kg = hs... hm = 2.393.0itling Szi Sp+a sig sRAL h - 897.719 + (0.001143)(120-19.08N)X100 he = 909.5567 kJ / kg Apply Stedy Hlow Egery Ezuation for closed feed water heatPaul work we = (höchs) xmg = 0.6054 MW Heat supplied Q. = (h-hy) + (1-4) (hy ha) 3244.910 Kolly 2017 = 162.2.45 MW Thermal Ef